OpenStudy (anonymous):

x^2 + 2x + y^2 + 4y = 9. What is the center?

OpenStudy (anonymous):

An explanation of how to find it would also be very helpful because I have another like this. Thank you. (:

OpenStudy (anonymous):

Sorry I thought it was Radius Center = (-g,-f)

OpenStudy (anonymous):

standard equation is \[x^2 + y^2 + 2gx + 2fy +c = 0 \]

OpenStudy (anonymous):

Now when we compare coefficient g=1 and f =2

OpenStudy (anonymous):

so that makes center \[(-1,-2)\]

myininaya (myininaya):

x^2+2x+1+y^2+4y+4=9+1+4 (x+1)^2+(y+2)^2=14 center is (-1,-2) gj ishaan